Find the volume generated by rotating the region bounded by y = x, x = 1, and y2 = 4x, about the x-axis.

1
Expert's answer
2021-06-12T05:09:46-0400

Given, the region is bounded by "y=x, y^2=4x" and "x=1" .

Volume generated is given by,

"v=\\pi\\int_a^b ([f(x)]^ 2\u2212[g(x)] ^2)dx\\\\\n=\\pi\\int_0^1((2\\sqrt x)^ 2\u2212x^ 2 )dx\\\\\n=\\pi\\int_0^1(4x\u2212x ^2)\\\\\n=\\pi(2x ^2\u2212 \\frac{x^3}{3})_0^1 \\\\\n=\\pi(2\u2212 \\frac{1}{3})\\\\\n= \\frac{5\\pi}{3}\n\u200b"

Thus, volume is "\frac{5\pi}{3}" .